### v1

@@ Abstract
End-to-end learning has shown great potential in autonomous parking, yet the lack of publicly available datasets limits reproducibility and benchmarking. While prior work introduced a visual-based parking model and a pipeline for data generation, training and close-loop test, the dataset itself was not released. To bridge this gap, we create and open-source a high-quality dataset for end-to-end autonomous parking. Using the original model, we achieve an overall success rate of 85.16% with lower average position and orientation errors (0.24 meters and 0.34 degrees). The dataset can be found in https://pan.baidu.com/s/16LSSfF_Qar3e-s_dseDt1g?pwd=wybs.
keywords:
@@
Autonomous Driving, Autonomous Parking, End-to-End Neural Network, Open Dataset, Deep Learning
1 Introduction
@@
The field of autonomous driving has advanced significantly, with a focus on improving autonomous parking. Traditional multi-stage pipelines, including perception, mapping, planning, and control, often suffer from sensor noise, model uncertainties, and rigid handcrafted rules, leading to error accumulation and poor adaptability in complex environments.
@@
End-to-end learning has emerged as a potential alternative by directly mapping sensor inputs to control actions, streamlining the pipeline, reducing reliance on handcrafted features, and improving adaptability. This approach has shown success in lane keeping, object detection, and collision avoidance. Inspired by this, previous work proposed an end-to-end parking method integrating visual inputs and motion states to generate control signals.
@@
However, a key challenge in this domain is the lack of publicly available datasets for training and evaluation. While Yunfan Yang et al.[1] introduced a model and shared a pipeline ranging from data generation to training and closed-loop test, their dataset remains unreleased. To address this gap, we create and open-source an end-to-end autonomous parking dataset based on their environmental setup.
@@
In this paper, we present the iterative process of creating a parking dataset that captures a wide variety of parking scenarios. Using the model architecture proposed by the original authors, we trained the model on the dataset and achieved an overall success rate of 85.16%, which is comparable to the original authors’ result of 91.41%. Additionally, we observed significant reductions in both average position and orientation errors (0.24 meters and 0.34 degrees) compared to the baseline (0.30 meters and 0.87 degrees), indicating improved parking precision.
@@
2 Related Works
In this part, the development of Transformer model and BEV model will be reviewed, and based on this, the application of these algorithms in the field of automatic parking will be introduced.
2.1 Transformer model
Transformer models have revolutionized AI since their introduction in the paper “Attention Is All You Need” by Vaswani et al. in 2017 [2]. It was first applied to Natural Language Processing (NLP) where it achieved significant success, such as BERT[3], GPT[4] and T5[5]. In the field of computer vision (CV), Vision Transformers (ViTs)[6] are able to outperform CNNs in image classification tasks by capturing both global and local features and understanding image-wide context. DETR[7] is a successful end-to-end object detection framework with a Transformer encoder-decoder architecture.
@@ 2.1 Transformer model
2.2 BEV representation
The Bird’s Eye View (BEV) is widely used in autonomous driving for tasks like semantic segmentation and object detection, providing a top-down perspective of the environment. Since 2D image features lack depth, LSS[16] method learns depth distributions and uses cameras’ intrinsic and extrinsic matrices to transform the frustum into BEV. Based on LSS, BEVDepth[17] improves depth estimation with explicit supervision. Transformer models, as in Transfuser[8], are used for BEV generation by fusing LiDAR and image features. ST-P3[18] is an end-to-end vision-based autonomous driving framework that leverages spatial-temporal BEV feature learning for perception, prediction and planning.
2.3 Autonomous driving datasets
@@
3 Method
The innovative architecture proposed by the original authors is continued in this work, where a neural network maps inputs to control signal outputs.
3.1 Problem Definition
@@ 3.2 Inputs and Outputs
Outputs: The model predicts control signals for the next four steps with 0.1s interval. The outputs include normalized acceleration ($\text{acc}_{t}\in[-1,1]$), steering angle ($\text{steer}_{t}\in[-1,1]$), and gear selection ($\text{gear}_{t}\in\{0,1\}$). Inspired by Pix2Seq [25], control prediction is framed as a language modeling task, where signals are tokenized and predicted sequentially. Acceleration and steering are discretized into 201 values $[0,200]$, with 0 indicating full brake/left and 200 indicating full throttle/right. For gear state, 0 stands for forward gear and 200 means reverse gear. The sequence format is:
@@ 3.4 Loss
4 Dataset Generation
4.1 Implementation details
The simulator is CARLA 0.9.11, which keeps the same as the one in the original paper. The Figure 1 shows the distribution of slots in the parking lot, where the red letter “T” designates the target slot. The red parking slots 2-2, …, 2-16, 3-2, …, 3-16 which end with even number are used for data generation and the brown parking slots 2-1, …, 2-15, 3-1, …, 3-15 which end with odd number are used for closed-loop test. Most of the settings are the same as the original ones. The ego vehicle is initialized at a random position of the path between the second and the third row, heading parallel with the path.
@@ 4.2.1 Keyboard Control
4.2.2 Insufficient scenes and dataset size
@@
4.2.3 Lack of parking slot with shadow of streetlight
@@
4.2.4 Trouble of parking slots near the edge of the parking lot
@@
4.3 Dataset overview
Each route data of the dataset was generated via the script provided by original authors. The sampling frequency is 10 Hz. The data for model input is RGB images from 4 cameras, extrinsic and intrinsic matrix for each camera, target point and motion state of ego vehicle, while control values, depth and segmentation are ground-truth labels. The range and resolution of the control values are listed in Table 1. 5 versions of dataset were created for comparison experiments by one expert driver: Gen 1A, 1B, 1C, 2A, 2B.
Table 1: Range and resolution of control values
@@ 4.3 Dataset overview
Reverse {0, 1} /
[…]
In Gen 1, after driving straight, the vehicle stops at a position slightly farther from the target parking slot. It then gradually increases the steering angle while reversing. When both rear wheels cross the extended boundary line of the target slot, the steering angle is increased to its maximum. Subsequently, the steering and throttle values are gradually reduced. As the vehicle’s front approaches alignment, the steering angle is fine-tuned. Once aligned, the throttle is further adjusted in reverse until the vehicle reaches the desired parking position, at which point the brake is fully applied. During the model test for parking slot 2-1 and 3-1, the initialization of the vehicle is outside the parking lot, with half of the body on the road. Therefore, during the data generation for 2-2 and 3-2, original initialization was modified to make the vehicle generated at that position in 10 routes, so that the model would have potential to handle this type of situation.
In Gen 2, the vehicle stops closer to the target parking slot after driving straight compared to Gen 1. It then reverses with the steering angle at its maximum. As the rear approaches the slot, the steering and throttle values are gradually decreased. The subsequent operations are the same as those in Gen 1. There is a slight change for initialization compared to Gen 1, in order to remove the instability caused by roadside shoulder. In Gen 2, the initial positions of parking slots 2-1 and 3-1 during testing are always within the parking lot, whereas for some routes of slots 2-2 and 3-2 during data generation, the vehicle is initialized near the edge of the parking lot to ensure coverage of edge-adjacent scenarios.
The selection of random seeds for each dataset is listed in Table 2. The random seed ranges refer to the seeds employed in the scene generation process, which can change the total number of vehicles and the type of vehicle in each parking slot.
Table 2: Number of trajectories in each dataset with different random seeds
@@
Seed 16 to 143 96 / 32  [^1] Gen 2A has more scenes. For example, random seeds 16, 32, 48, 64, 80, 96, 112 and 128 are assigned to the first parking lot for generation.
Slot 3-8 with shadow, seed 11, 27 12 / 4  [^2] Additionally, routes are generated for slot 3-8 and 3-10 with shadow separately after modifying the sun azimuth angle and altitude angle. For example, when random seed 11 is assigned to slot 3-8 in scene generation, 6 routes are for training set and 2 routes are for validation set. The same for seed 27. 12 / 4
Slot 3-10 with shadow, seed 12, 28 12 / 4  [^2] Additionally, routes are generated for slot 3-8 and 3-10 with shadow separately after modifying the sun azimuth angle and altitude angle. For example, when random seed 11 is assigned to slot 3-8 in scene generation, 6 routes are for training set and 2 routes are for validation set. The same for seed 27. 12 / 4
Slot 3-8 with shadow, seed 91, 107, 123, 139 (3 / 1)  [^3] For slot 3-8 and 3-10 in Gen 2A, the routes with shadow in the slots are not generated separately, but are included in the cases with random seeds ranging from 16 to 143 mentioned above. Therefore, these routes are not counted in the total.
Slot 3-10 with shadow, seed 28, 44, 60, 76 (3 / 1)  [^3] For slot 3-8 and 3-10 in Gen 2A, the routes with shadow in the slots are not generated separately, but are included in the cases with random seeds ranging from 16 to 143 mentioned above. Therefore, these routes are not counted in the total. (1 / 1)  [^4] The routes for training and validation are assigned seeds 28 and 44, respectively, both of which are included in the cases mentioned above.
@@ 5.1.1 Iteration of datasets
Similar to Gen 1A, 128 trajectories were generated for Gen 2A using random seeds 0 to 15. The difference is that the straight-line distance before reversing was shortened to prevent the vehicle from driving out of the parking lot. To enhance the diversity, more random seeds are used to generate scenes for the extra 128 routes, as Table 3 illustrates.
Table 3: Random seed and initial position assignment w.r.t weather settings in training and validation sets in Gen 2A
@@
[…]
00footnotetext: Each random seed is assigned to one of 16 parking slots with different weather settings to ensure a diverse dataset. Weather 0 is default weather, in which the shadow of streetlight is projected on slot 3-9. In Weather 1, the sun angles are set to project the shadow on slot 3-10, while in Weather 2, the shadow is on slot 3-8.
@@ 5.3 Results
5.3.1 Comparison of Gen 1A, 1B, and 1C
The model trained on the Gen 1A dataset performed poorly in testing, achieving a TSR of only 44.792% after 155 epochs of training. For comparison, the model trained on the Gen 1B dataset achieved a TSR of 65.625% in testing after 155 epochs with the same learning rate of $1\times 10^{-4}$, benefiting from the doubled dataset size and increased scenario diversity.
Table 4: Gen 1A’s Performance
@@ 5.3.1 Comparison of Gen 1A, 1B, and 1C
[…]
@@
Table 5: Comparison of GEN 1B’s Performance under Different Initial Learning Rates
@@
[…]
As Table 6 demonstrates, with the additional parking trajectories in shadowed slots, Gen 1C achieved a higher TSR in shadowed parking slots and outperformed Gen 1B in overall TSR. Specifically, with an initial learning rate of $1\times 10^{-4}$, the model trained on Gen 1C for 150 epochs achieved a TSR 7.25 percentage points higher than the model trained on Gen 1B for 155 epochs. Similarly, with an initial learning rate of $7.5\times 10^{-5}$, the model trained on Gen 1C for 155 epochs outperformed the Gen 1B model trained for the same number of epochs by 11.46 percentage points.

### v2

@@ Abstract
While traditional autonomous driving methods with multi-stage pipelines suffer from lengthy processes, error accumulations and maintenance difficulties, the end-to-end method is designed to map the data of multiple sensors directly into motion control commands, with high flexibility, efficiency and generalization. Therefore, the end-to-end model has shown great potential in autonomous driving. Due to the low speed, low risk, and low complexity characteristics of autonomous parking scenarios, end-to-end methods can be applied to autonomous parking systems earlier. However, the lack of publicly available datasets limits reproducibility and benchmarking in the early research stage of end-to-end autonomous parking. While prior work introduced a visual-based parking model and a pipeline for data generation, training and closed-loop test, the dataset itself was not released. To bridge this gap, we work on creating large end-to-end autonomous parking datasets in CARLA based on the prior work ‘E2E Parking’. Keyboard control is replaced by Handle Controller to improve usability, efficiency, and operational precision. During the iterative process of dataset generation, we evaluate the effect of different factors on the parking performance of the controlled vehicle, including diverse scenes generated by multiple random seeds, the position of the roadside object’s shadow dependent on weather setting, dataset size, initial learning rate and training epochs. We recommend generating at least 2 scenes for each parking slot with different random seeds, where 8 trajectories with different initial positions are collected for each scene. Weather settings should be modified to make the dataset include scenes with shadow projected on the target slot. Experiments demonstrate that an initial learning rate of $7.5\times 10^{-5}$ and epochs set as around 150 can make the model perform better. After several iterations, we are able to open-source a high-quality dataset for end-to-end autonomous parking. Using the original transformer-based model, we achieve an overall success rate of 85.16% with lower average position and orientation errors (0.24 meters and 0.34 degrees). The dataset can be found in https://github.com/KejiaGao/e2e-parking-carla-dataset.
Index Terms:
@@
Autonomous Driving, Autonomous Parking, End-to-End Neural Network, Imitation Learning, Transformer.
I Introduction
@@
The field of autonomous driving has advanced significantly, with a focus on improving autonomous parking. Traditional multi-stage pipelines, including environment perception, mapping, path planning, and motion control, often suffer from sensor noise, model uncertainties, and rigid handcrafted rules, leading to error accumulation and poor adaptability in complex environments.
@@
End-to-end method has emerged as a potential alternative by directly mapping sensor data to motion control commands, streamlining the pipeline, reducing reliance on handcrafted features, and improving adaptability. This approach has shown success in lane keeping, object detection, and collision avoidance. Inspired by this, previous work proposed an end-to-end parking method integrating visual inputs and motion states to generate control commands.
@@
However, a key challenge in this domain is the lack of publicly available datasets for training and evaluation. While Yang et al. [1] introduced a model and shared a pipeline ranging from data generation to training and closed-loop test, their dataset remains unreleased and the pre-trained model they have published has only an overall success rate of around 75%. To address this gap, we create and open-source an end-to-end autonomous parking dataset based on their environmental setup.
@@
In this paper, we present the iterative process of creating a parking dataset that captures a wide variety of parking scenarios. Using the model architecture proposed by the original authors, we trained the model on the dataset and achieved an overall success rate of 85.16%, which can replicate and exceed the performance of the released pre-trained model. Furthermore, it is comparable to the original authors’ result of 91.41% from their unpublished model. Additionally, we observed significant reductions in both average position and orientation errors (0.24 meters and 0.34 degrees) compared to the baseline (0.30 meters and 0.87 degrees), indicating improved parking precision.
@@
II Related Works
In this part, the development of Transformer model and BEV model used in autonomous driving will be reviewed, and based on this, the application of these algorithms in the field of automatic parking will be introduced.
II-A Transformer Model
Transformer models have revolutionized AI since their introduction in the paper “Attention Is All You Need” [2] by Vaswani et al. in 2017. It was first applied to Natural Language Processing (NLP) where it achieved significant success, such as BERT [3], GPT [4] and T5 [5]. In the field of computer vision (CV), Vision Transformers (ViTs) [6] are able to outperform CNNs in image classification tasks by capturing both global and local features and understanding image-wide context. DETR [7] is a successful end-to-end object detection framework with a Transformer encoder-decoder architecture.
@@ II-A Transformer Model
II-B BEV Representation
The Bird’s-Eye View (BEV) is widely used in autonomous driving for tasks like semantic segmentation and object detection, providing a top-down perspective of the environment. Since 2D image features lack depth, LSS [16] method learns depth distributions and uses cameras’ intrinsic and extrinsic matrices to transform the frustum into BEV. Based on LSS, BEVDepth [17] improves depth estimation with explicit supervision. Transformer models, as in Transfuser [8], are used for BEV generation by fusing LiDAR and image features. ST-P3 [18] is an end-to-end vision-based autonomous driving framework that leverages spatial-temporal BEV feature learning for perception, prediction and planning.
II-C Autonomous Driving Datasets
@@
III Method
The innovative architecture [1] proposed by the original authors is continued in this work, where a neural network maps inputs to control command outputs.
III-A Problem Definition
@@ III-B Inputs and Outputs
Outputs: The model predicts control signals for the next four steps with 0.1s interval. The outputs include normalized acceleration ($\text{acc}_{t}\in[-1,1]$), steering angle ($\text{steer}_{t}\in[-1,1]$), and gear selection ($\text{gear}_{t}\in\{0,1\}$). Inspired by Pix2Seq [25], control prediction is framed as a language modeling task, where signals are tokenized and predicted sequentially. Acceleration and steering are discretized into 201 values [0,200][0,200][ 0 , 200 ], with 0 indicating full brake/left and 200 indicating full throttle/right. For gear state, 0 stands for forward gear and 200 means reverse gear. The sequence format is:
@@ III-D Loss
IV Dataset Generation
IV-A Implementation Details
The simulator is CARLA 0.9.11, which keeps the same as the one in the original paper. Figure 1 shows the distribution of slots in the parking lot, where the red letter “T” designates the target slot. The red parking slots 2-2, …, 2-16, 3-2, …, 3-16 which end with even number are used for data generation and the brown parking slots 2-1, …, 2-15, 3-1, …, 3-15 which end with odd number are used for closed-loop test. Most of the settings are the same as the original ones. The ego vehicle is initialized at a random position of the path between the second and the third row, heading parallel with the path.
@@ IV-B1 Keyboard Control
IV-B2 Insufficient scenes and dataset size
@@
IV-B3 Lack of parking slot with shadow of streetlight
@@
IV-B4 Trouble of parking slots near the edge of the parking lot
@@
IV-C Dataset Overview
Each route data of the dataset was generated via the script provided by original authors. The sampling frequency is 10 Hz. The data for model input is RGB images from 4 cameras, extrinsic and intrinsic matrix for each camera, target point and motion state of ego vehicle, while control values, depth and segmentation are ground-truth labels. The range and resolution of the control values are listed in Table I. 5 versions of dataset were created for comparison experiments by one expert driver: Gen 1A, 1B, 1C, 2A, 2B.
TABLE I: Range and resolution of control values
@@ IV-C Dataset Overview
Reverse {0, 1}
In Gen 1, after driving straight, the vehicle stops at a position slightly farther from the target parking slot, as is shown in Figure 2(b). It then gradually increases the steering angle while reversing. When both rear wheels cross the extended boundary line of the target slot (like Figure 2(d)), the steering angle is increased to its maximum. Subsequently, the steering and throttle values are gradually reduced. As the vehicle’s front approaches alignment, the steering angle is fine-tuned (like Figure 2(e)). Once aligned, the throttle is further adjusted in reverse until the vehicle reaches the desired parking position (Figure 2(f)), at which point the brake is fully applied. During the model test for parking slot 2-1 and 3-1, the initialization of the vehicle is outside the parking lot, with half of the body on the road. Therefore, during the data generation for 2-2 and 3-2, original initialization was modified to make the vehicle generated at that position in 10 routes, so that the model would have potential to handle this type of situation.
(a) Initialization
(b) Stop
(c) Reverse and steer
(d) Reverse and steer
(e) Fine tuning
(f) End
Figure 2: Parking strategy in Gen 1
In Gen 2, the vehicle stops closer to the target parking slot after driving straight compared to Gen 1, as Figure 3(b) shows. It then reverses with the steering angle at its maximum. As the rear approaches the slot, the steering and throttle values are gradually decreased (like Figure 3(c) and 3(d)). The subsequent operations are the same as those in Gen 1. There is a slight change for initialization compared to Gen 1, in order to remove the instability caused by roadside shoulder. In Gen 2, the initial positions of parking slots 2-1 and 3-1 during testing are always within the parking lot, whereas for some routes of slots 2-2 and 3-2 during data generation, the vehicle is initialized near the edge of the parking lot to ensure coverage of edge-adjacent scenarios.
(a) Initialization
(b) Stop
(c) Reverse and steer
(d) Reverse and steer
(e) Fine tuning
(f) End
Figure 3: Parking strategy in Gen 2
The selection of random seeds for each dataset is listed in II. The random seed ranges refer to the seeds employed in the scene generation process, which can change the total number of vehicles and the type of vehicle in each parking slot.
V Experiments
TABLE II: Number of trajectories in each dataset with different random seeds
@@ V Experiments
Seed 16 to 143 96 / 32111Gen 2A has more scenes. For example, random seeds 16, 32, 48, 64, 80, 96, 112 and 128 are assigned to the first parking lot for generation.
3-8 (shadow), seed 11, 27 12 / 4 12 / 4
3-10 (shadow), seed 12, 28 12 / 4222Additionally, in Gen 1C, routes are generated for slot 3-8 and 3-10 with shadow separately after modifying the sun azimuth angle and altitude angle. For example, when random seed 11 is assigned to slot 3-8 in scene generation, 6 routes are for training set and 2 routes are for validation set. The same for seed 27. 12 / 4
3-8 (shadow), seed 91, 107, 123, 139 (3 / 1)
3-10 (shadow), seed 28, 44, 60, 76 (3 / 1) [^3] For slot 3-8 and 3-10 in Gen 2A, the routes with shadow in the slots are not generated separately, but are included in the cases with random seeds ranging from 16 to 143 mentioned above. Therefore, these routes are not counted in the total. (1 / 1) [^4] The routes for training and validation are assigned seeds 28 and 44, respectively, both of which are included in the cases mentioned above.
@@ V-A1 Iteration of datasets
Similar to Gen 1A, 128 trajectories were generated for Gen 2A using random seeds 0 to 15. The difference is that the straight-line distance before reversing was shortened to prevent the vehicle from driving out of the parking lot. To enhance the diversity, more random seeds are used to generate scenes for the extra 128 routes, as III illustrates. Each random seed is assigned to one of 16 parking slots with different weather settings to ensure a diverse dataset. Weather 0 is default weather, in which the shadow of streetlight is projected on slot 3-9. In Weather 1, the sun angles are set to project the shadow on slot 3-10, while in Weather 2, the shadow is on slot 3-8.
TABLE III: Random seed and initial position assignment w.r.t weather settings in training and validation sets in Gen 2A
@@ V-C Results
V-C1 Details of Gen 1A testing
According to Table IV, the model trained on the Gen 1A dataset performed poorly in testing, achieving a TSR of only 44.792% after 155 epochs of training. For comparison, the model trained on the Gen 1B dataset achieved a TSR of 65.625% in testing after 155 epochs with the same learning rate of $1\times 10^{-4}$, benefiting from the doubled dataset size and increased scenario diversity.
TABLE IV: Gen 1A’s Performance
@@ V-C1 Details of Gen 1A testing
Table V presents the test results of the best model trained on Gen 1A. Each of the 16 parking slot is tested for 6 times with different initial positions. For parking slots 2-1, 2-3, 3-1, and 3-9, the model completely failed to park the vehicle into the target slot. For slots 2-1 and 3-1, the main problem is the high OR, which means the vehicle frequently exits from the parking lot abnormally. For slot 2-3, the vehicle collides with the neighboring vehicle very often. For slot 3-9, the model has a misunderstanding with the shadow on the ground and refuses to park the vehicle into the target slot, which causes timeout and collision. For other parking slots, the model can’t ensure a robust performance and the TSR varies from slot to slot.
TABLE V: Closed-loop results including 96 test cases for the best model trained on Gen 1A
TaskIdx TSR (%) TFR (%) NTSR (%) NTFR (%) CR (%) OR (%) TR (%) APE (m) AOE (deg) APT (s) AIT (s)
2-1 0.000 0.000 16.667 0.000 16.667 50.000 16.667 0.080
2-3 0.000 0.000 16.667 0.000 66.667 0.000 16.667 0.082
2-5 50.000 0.000 33.333 0.000 0.000 0.000 16.667 0.483 0.696 21.667 0.081
2-7 66.667 0.000 0.000 0.000 33.333 0.000 0.000 0.486 0.686 22.000 0.082
2-9 83.333 0.000 0.000 0.000 0.000 0.000 16.667 0.555 1.154 23.873 0.081
2-11 50.000 0.000 0.000 0.000 50.000 0.000 0.000 0.566 1.868 22.056 0.084
2-13 50.000 16.667 0.000 0.000 33.333 0.000 0.000 0.476 0.921 22.289 0.087
2-15 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.387 0.951 22.178 0.079
3-1 0.000 0.000 0.000 0.000 16.667 66.667 16.667 0.079
3-3 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.387 1.244 23.961 0.080
3-5 33.333 0.000 0.000 0.000 0.000 50.000 16.667 0.307 1.087 24.550 0.080
3-7 66.667 0.000 16.667 0.000 16.667 0.000 0.000 0.446 1.515 21.775 0.080
3-9 0.000 0.000 0.000 0.000 33.333 0.000 66.667 0.080
3-11 50.000 0.000 16.667 0.000 0.000 33.333 0.000 0.445 1.738 22.456 0.079
3-13 16.667 0.000 0.000 0.000 0.000 33.333 50.000 0.348 0.903 24.233 0.079
3-15 50.000 0.000 0.000 0.000 0.000 16.667 33.333 0.285 0.279 24.133 0.079
Avg 44.792 1.042 6.250 0.000 16.667 15.625 15.625 0.431 1.087 22.931 0.081
V-C2 Details of Gen 1B testing
@@
TABLE VI: Comparison of GEN 1B’s Performance under Different Initial Learning Rates
@@
Table VII depicts the detailed test results of the best model trained on Gen 1B. With the help of a larger dataset, the model demonstrates a partial ability to handle parking tasks at target slots located at the edge of the parking area. For slot 3-9 with shadow on the ground, the model still can’t deal with the task at all. However, the TSRs for other parking slots have improved to varying degrees, with the overall average TSR increasing by 27.083 percentage points compared to Gen 1A, demonstrating the great potential of the larger size and more scenes of dataset.
TABLE VII: Closed-loop results including 96 test cases for the best model trained on Gen 1B
TaskIdx TSR (%) TFR (%) NTSR (%) NTFR (%) CR (%) OR (%) TR (%) APE (m) AOE (deg) APT (s) AIT (s)
2-1 33.333 0.000 0.000 0.000 0.000 66.667 0.000 0.409 0.295 23.100 0.078
2-3 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.492 0.685 22.150 0.076
2-5 83.333 16.667 0.000 0.000 0.000 0.000 0.000 0.415 0.446 22.313 0.078
2-7 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.484 0.276 23.433 0.080
2-9 83.333 0.000 0.000 0.000 0.000 0.000 16.667 0.445 0.493 23.653 0.078
2-11 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.458 0.718 21.239 0.080
2-13 66.667 0.000 33.333 0.000 0.000 0.000 0.000 0.436 0.231 22.983 0.114
2-15 83.333 0.000 16.667 0.000 0.000 0.000 0.000 0.306 0.314 22.927 0.077
3-1 16.667 0.000 0.000 0.000 0.000 83.333 0.000 0.217 0.270 24.600 0.076
3-3 83.333 0.000 0.000 0.000 16.667 0.000 0.000 0.245 0.389 22.047 0.077
3-5 83.333 0.000 0.000 0.000 0.000 0.000 16.667 0.296 0.630 24.007 0.077
3-7 83.333 0.000 0.000 0.000 0.000 0.000 16.667 0.327 0.753 22.527 0.077
3-9 0.000 0.000 33.333 0.000 0.000 0.000 66.667 0.077
3-11 66.667 0.000 0.000 0.000 0.000 0.000 33.333 0.405 0.447 24.517 0.077
3-13 66.667 0.000 0.000 0.000 0.000 0.000 33.333 0.181 0.248 25.217 0.077
3-15 100.000 0.000 0.000 0.000 0.000 0.000 0.000 0.209 0.396 23.206 0.077
Avg 71.875 1.042 5.208 0.000 1.042 9.375 11.458 0.355 0.439 23.195 0.080
V-C3 Details of Gen 1C testing
As Table XIII demonstrates, with the additional parking trajectories in shadowed slots, Gen 1C achieved a higher TSR in shadowed parking slots and outperformed Gen 1B in overall TSR. Specifically, with an initial learning rate of $1\times 10^{-4}$, the model trained on Gen 1C for 150 epochs achieved a TSR 7.25 percentage points higher than the model trained on Gen 1B for 155 epochs. Similarly, with an initial learning rate of $7.5\times 10^{-5}$, the model trained on Gen 1C for 155 epochs outperformed the Gen 1B model trained for the same number of epochs by 11.46 percentage points.